OpenStudy (anonymous):

26 cm is how many mm? 10 cm is how many mm? 300 cm is how many meters? 130 cm how tall is that in meters? 130 mm is how many cm? 200 cm is how many meters? 100 cm is how many meters? 70 mm is how many cm? 188 cm is how many mm? have to show how to get answers. how to do the problems.

OpenStudy (anonymous):

1 cm = 10 mm so 26 cm = 26*10 mm = 260 mm

OpenStudy (anonymous):

1 cm = 1/100 m so 300 cm = 300/100 m = 3 m

OpenStudy (anonymous):

26 centimeters = 260 millimeters 10 centimeters = 100 millimeters 300 centimeters = 3 meters

OpenStudy (anonymous):

what does * mean?

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

* stands for multiplication

OpenStudy (anonymous):

130 millimeters = 13 centimeters 100 centimeters = 1 meter 200 centimeters = 2 meters 70 millimeters = 7 centimeters 188 centimeters = 1880 millimeters

OpenStudy (anonymous):

Sorry I didn't explain there was a lot of question to answers but hope this helps!:)

OpenStudy (anonymous):

it is ok i am writing it down so that i can understand it thank you very much